1.how to find oxidation number?
2.how to give IUPAC nomenclature?
3.what are the rules for iupac naming?
 

Grade:11

1 Answers

Arun
5 years ago

There are several rules to find the oxidation state of an element present in a compound or ion. Some of them are given below.

  1. The oxidation state of elements which occur in their naturally existing forms are considered to be zero. Ex: H2, O2, F2, Mg, etc. ( This is the general way they are found in the world so they are considered as zero.
  2. In monoatomic ions the charge of the ion is considered as their oxidation number. Ex: Mg2+= +2, Fe3+= +3
  3. The oxidation state of the elements in Group I and Group II when existing as compounds is always respectively equal to +1 and +2. Ex: Na2CO3; Na= +1, KCl; K= +1, SrCr2O7; Sr= +2
  4. The oxidation state of Fluorine in any combined state except F2 is always -1. ( Fluorine never has positive oxidation states.)
  5. In polyatomic ions the summation of the oxidation states of all species present in the ion is equal to the overall charge of the ion. Ex: MnO4-; Charge of the ion= -1, Oxidation of Manganese= +7, Oxidation of Oxygen= -2(There are four oxygens so total is -8) Then +7 + (-8)= -1
  6. The summation of oxidation states of all species present in a neutral compound is equal to zero. Ex: NaS; Na= +1( Two sodiums so total is +2), S= -2 Then +2 + (-2)=0
  7. Oxygen mostly has the -2 oxidation in compounds except when it is bonded to Fluorine(which is more electronegative than oxygen) or is part of the peroxide or superoxide ion where it is -1 and -1/2 respectively. Ex: Na2O( Sodium oxide); Na= +1( Rule 3) Considering +2( Two Na) + X= 0(Rule 6), X= Oxidation state of Oxygen= -2

NaO2( Sodium peroxide); Na= +1 (Rule 3) using (Rule 6) = +2 + 2X = 0, X= -1

BaO2( Barium peroxide); Ba= +2 (Rule 3) using (Rule 6) = +2 + 2X=0, X= -1

CsO2( Caesium superoxide); Cs= +1( Rule 3) using (Rule 6) = +1 + 2X=0, X= -1/2

8. Certain p-block elements and d-block and f-block elements can show varied oxidation states in different compounds.Ex: p-block elements= Sulphur can show oxidation states -2, -1, +4, +6 ( Na2S, Na2S2, SO2, H2SO4)

d-block elements= Manganese= +2,+4,+7 ( MnSO4, MnO2, KMnO4) Iron = +2, +3 ( FeCl2, FeCl3)